OpenStudy (anonymous):

Find the focus of the parabola x = –2(y – 3)^2 - 2.

OpenStudy (campbell_st):

re write the equation x + 2 =-2(y-3)^2 now divide both sides by -2 \[\frac{-1}{2} (x +2) = (y -3)^2\] the general from is 4a(x -h) = (y -k)^2 (h, k) is the vertex and a is the focal length... the focus will be (h - a, k) the parabola is concave left...